# Zhao* Vu* Wang

### 05 Mar 2026 v1

@@ Abstract
Neural codec language models enable high-quality discrete speech synthesis, yet their inference remains vulnerable to token-level artifacts and distributional drift that degrade perceptual realism. Rather than relying on preference optimization or retraining, we propose MSpoof-TTS, a training-free inference framework that improves zero-shot synthesis through multi-resolution spoof guidance. We introduce a Multi-Resolution Token-based Spoof Detection framework that evaluates codec sequences at different temporal granularities to detect locally inconsistent or unnatural patterns. We then integrate the spoof detectors into a hierarchical decoding strategy, progressively pruning low-quality candidates and re-ranking hypotheses. This discriminator-guided generation enhances robustness without modifying model parameters. Experiments validate the effectiveness of our framework for robust and high-quality codec-based speech generation.
@@ 1 Introduction
Neural codec language models have recently become a practical and effective approach for zero-shot speech synthesis [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. By modeling speech as sequences of discrete codec tokens with autoregressive or transformer architectures, these systems streamline the synthesis pipeline and naturally adopt scalable decoding strategies from large language models. In practice, however, generation in the discrete token space can be fragile. Small inconsistencies at the token level may accumulate during autoregressive decoding, resulting in audible artifacts, locally unnatural transitions, or a gradual drift away from natural speech characteristics [11]. Since next-token prediction objectives do not explicitly constrain such behavior, these issues are often difficult to detect or correct during inference.
@@
To mitigate these decoding instabilities, existing methods generally fall into two categories. One line of work introduces additional supervision or reward-driven objectives to better align model outputs with perceptual or distributional goals [11, 12, 13, 14, 15]. For example, SpeechAlign [11] reduces the mismatch between golden and synthesized codec tokens through preference-based optimization. Building on similar ideas, [12] leverages human feedback signals as training supervision to improve perceptual quality, while [13] integrates differentiable reward signals directly into the optimization objective to refine generation behavior. In addition, [14] introduces an auxiliary reverse inference objective to enhance robustness under distributional shift. Although effective, these approaches often require retraining, iterative optimization, or carefully curated data, which increases computational cost and system complexity.
@@
Another line of work focuses on decoding-time adjustments, including repetition control, alignment constraints, or modified sampling strategies [16, 7, 17]. VALL-E 2 [16] introduces repetition-aware sampling to mitigate degeneration during autoregressive decoding. ELLA-V [7] improves stability by refining decoding strategies under long-context and zero-shot settings. VALL-E R [17] enhances robustness through inference-time alignment constraints and decoding refinements. These methods are simple to apply and avoid retraining, but they primarily target specific failure patterns rather than explicitly assessing whether the generated token sequence remains globally consistent or locally natural.
@@
A related idea has been explored in controllable text generation, where external classifiers are incorporated directly during decoding to guide model outputs without retraining the base language model [18, 19]. Instead of optimizing the generator through reinforcement learning, these approaches influence generation at inference time by adjusting token probabilities or re-ranking candidate sequences based on auxiliary evaluation signals. Such strategies demonstrate that generation behavior can be steered through an explicit evaluation mechanism applied during decoding, rather than by modifying model parameters. Motivated by this perspective, we introduce a spoof detection model as an authenticity evaluator and integrate it into the decoding process to guide discrete speech generation.
@@
To this end, spoofed and deepfake speech detection provides a natural foundation. Detection of synthetic or manipulated speech has been extensively studied in the continuous audio domain [20, 21, 22, 23, 24], with large-scale benchmarks such as the ASVspoof challenge series [25, 26, 27, 28] driving substantial progress in waveform-level countermeasure design. More recently, codec-based deepfake datasets have revealed additional challenges introduced by neural codec synthesis [29, 30, 31]. However, existing spoof detection systems are primarily formulated on reconstructed audio signals and are designed for post-hoc classification. They do not operate on discrete codec token sequences, nor are they intended to guide generation during decoding.
@@ 2.1 Overview
Pretrained codec-based language models have demonstrated strong capability in zero-shot speech synthesis by modeling discrete codec token sequences autoregressively. In this work, we adopt NeuTTS [^1] https://github.com/neuphonic/neutts , a pretrained codec-based TTS system, as the base generator, whose parameters remain fixed throughout our framework. While such models are trained to predict next tokens conditioned on ground-truth codec sequences, inference relies on previously generated tokens. This shift introduces a discrepancy between training and decoding, under which synthesized token sequences may progressively deviate from the natural codec distribution, as previously observed in [11].
@@ 2.2 Multi-Resolution Token-Level Spoof Detection
In addition, inspired by multi-scale discriminators in neural vocoders [32, 33], we introduce a resolution-based skip sampling strategy with downsampling rates $r\in\{1,2,5\}$. When $r=1$, the sequence remains at its native resolution, corresponding to the full 50-token span. Larger $r$ values subsample the sequence to form coarser token representations, encouraging the detector to uncover structural inconsistencies that may not be evident at the original granularity. While temporal cropping enhances coverage across time spans, skip sampling perturbs token resolution to probe structural regularity from a complementary perspective.
Each token segment is processed by an embedding layer followed by stacked Conformer [34] blocks to jointly capture local correlations and long-range dependencies. The resulting sequence representation is aggregated via adaptive pooling and passed to a lightweight classifier to predict the probability of the segment being real or synthetic. The detector is trained with the binary cross-entropy (BCE) objective and optimized independently of the TTS backbone. The architecture is shared across all resolution settings, while parameters are learned separately for each configuration. Specifically, we train five models: $\mathcal{M}_{50}$, $\mathcal{M}_{25}$, and $\mathcal{M}_{10}$ for contiguous cropping with segment lengths $L=50,25,10$, respectively, and $\mathcal{M}_{50\leftarrow 25}$ and $\mathcal{M}_{50\leftarrow 10}$ for skip-sampled variants obtained by downsampling 50-token segments with rates $r=2$ and $r=5$, respectively.
@@ 2.3 Hierarchical Spoof-Guided Sampling
To improve the decoding robustness of autoregressive codec language models, we propose a hierarchical spoof-guided sampling framework that integrates Entropy-Aware Sampling (EAS) with the multi-resolution token-level spoof detection introduced in Section 2.2. We first adopt Entropy-Aware Sampling (EAS), shown in Algorithm 1, as the base decoding strategy. EAS is adapted from repetition-aware sampling (RAS) proposed in VALL-E 2 [16], which penalizes recently generated tokens to mitigate looping behavior. However, conventional RAS relies on heuristic repetition counting and may over-suppress high-probability tokens, resulting in unstable decoding or excessive randomness. In contrast, EAS maintains a memory buffer that records competitive candidate tokens along with their rank positions and temporal age. The token-level penalty is modulated by inverse rank weighting and exponential temporal decay, with a clipping mechanism to prevent over-penalization. The adjusted distribution is then used for nucleus sampling, introducing entropy regularization while preserving distributional diversity.
@@ 3.1 Datasets
For spoof detection training, we use the LibriTTS [35] training split (approximately 100 hours of clean read English speech). LibriTTS is a multi-speaker corpus with careful segmentation and noise filtering. For each ground-truth utterance, we generate three synthetic counterparts using the same transcript but a different reference utterance from the same speaker under the default inference scheme. All real and synthetic waveforms are converted into discrete token sequences using the default speech tokenizer (NeuCodec).
For evaluation, experiments are conducted on both LibriSpeech [36] and LibriTTS (test split), which provide diverse speakers and clean TTS-style speech, respectively. These datasets enable evaluation under standard speech conditions using WER, speaker similarity, and perceptual quality metrics. To further assess robustness under challenging phonetic patterns, we additionally evaluate on the TwistList benchmark [37], which consists of linguistically constructed tongue twisters characterized by dense alliteration and repetitive phoneme patterns. We use the official test split and synthesize speech using the same conditioning and decoding pipeline as in the standard evaluation.
@@ 3.3 Evaluation Metrics
For objective evaluation, we report ASR-based intelligibility, speaker similarity, perceptual quality, and spoof detection performance. Intelligibility is measured using Word Error Rate (WER), computed with Whisper-large-v3 [38] [^2] https://huggingface.co/openai/whisper-large-v3 for all decoding schemes. Speaker similarity is evaluated using the WavLM [39] [^3] https://huggingface.co/microsoft/wavlm-base-plus-sv model by computing cosine similarity between L2-normalized embeddings of the generated utterance and another real utterance from the same speaker. Perceptual quality is assessed using neural MOS estimators, including NISQA [40] and MOSNet [41], both reflecting overall speech quality.

### 11 Apr 2026 v2

@@ Abstract
Neural codec language models enable high-quality discrete speech synthesis, yet their inference remains vulnerable to token-level artifacts and distributional drift that degrade perceptual realism. Rather than relying on preference optimization or retraining, we propose MSpoof-TTS, a training-free inference framework that improves zero-shot synthesis through multi-resolution spoof guidance. We introduce a Multi-Resolution Token-based Spoof Detection framework that evaluates codec sequences at different temporal granularities to detect locally inconsistent or unnatural patterns. We then integrate the spoof detectors into a hierarchical decoding strategy, progressively pruning low-quality candidates and re-ranking hypotheses. This discriminator-guided generation enhances robustness without modifying model parameters. Experiments validate the effectiveness of our framework for robust and high-quality codec-based speech generation. Audio samples are available at https://danny-nus.github.io/MSpoofTTS.github.io/.
@@ 1 Introduction
Neural codec language models have recently become a practical and effective approach for zero-shot speech synthesis [tan2024naturalspeech, ju2024naturalspeech, du2024cosyvoice, ye2025llasa, meng2025autoregressive, wangmaskgct, song2025ella, chen2025f5, zhao2025prosody, liang2026segment]. By modeling speech as sequences of discrete codec tokens with autoregressive or transformer architectures, these systems streamline the synthesis pipeline and naturally adopt scalable decoding strategies from large language models. In practice, however, generation in the discrete token space can be fragile. Small inconsistencies at the token level may accumulate during autoregressive decoding, resulting in audible artifacts, locally unnatural transitions, or a gradual drift away from natural speech characteristics [zhang2024speechalign]. Since next-token prediction objectives do not explicitly constrain such behavior, these issues are often difficult to detect or correct during inference.
@@
To mitigate these decoding instabilities, existing methods generally fall into two categories. One line of work introduces additional supervision or reward-driven objectives to better align model outputs with perceptual or distributional goals [zhang2024speechalign, hu2024robust, gao2025differentiable, chen2024enhancing, zhao2026comelsinger]. For example, SpeechAlign [zhang2024speechalign] reduces the mismatch between golden and synthesized codec tokens through preference-based optimization. Building on similar ideas, [hu2024robust] leverages human feedback signals as training supervision to improve perceptual quality, while [gao2025differentiable] integrates differentiable reward signals directly into the optimization objective to refine generation behavior. In addition, [chen2024enhancing] introduces an auxiliary reverse inference objective to enhance robustness under distributional shift. Although effective, these approaches often require retraining, iterative optimization, or carefully curated data, which increases computational cost and system complexity.
@@
Another line of work focuses on decoding-time adjustments, including repetition control, alignment constraints, or modified sampling strategies [chen2024vall, song2025ella, han2024vall]. VALL-E 2 [chen2024vall] introduces repetition-aware sampling to mitigate degeneration during autoregressive decoding. ELLA-V [song2025ella] improves stability by refining decoding strategies under long-context and zero-shot settings. VALL-E R [han2024vall] enhances robustness through inference-time alignment constraints and decoding refinements. These methods are simple to apply and avoid retraining, but they primarily target specific failure patterns rather than explicitly assessing whether the generated token sequence remains globally consistent or locally natural.
@@
A related idea has been explored in controllable text generation, where external classifiers are incorporated directly during decoding to guide model outputs without retraining the base language model [liu2021dexperts, dathathri2019plug]. Instead of optimizing the generator through reinforcement learning, these approaches influence generation at inference time by adjusting token probabilities or re-ranking candidate sequences based on auxiliary evaluation signals. Such strategies demonstrate that generation behavior can be steered through an explicit evaluation mechanism applied during decoding, rather than by modifying model parameters. Motivated by this perspective, we introduce a spoof detection model as an authenticity evaluator and integrate it into the decoding process to guide discrete speech generation.
@@
To this end, spoofed and deepfake speech detection provides a natural foundation. Detection of synthetic or manipulated speech has been extensively studied in the continuous audio domain [sanchez2015toward, zhang2021one, wang2024generalizable, sun2025contrastive, javed2022voice], with large-scale benchmarks such as the ASVspoof challenge series [wu2017asvspoof, todisco2019asvspoof, yamagishi2021asvspoof, wang2024asvspoof] driving substantial progress in waveform-level countermeasure design. More recently, codec-based deepfake datasets have revealed additional challenges introduced by neural codec synthesis [wu2024codecfake, chen2025codecfake+, du2025codecfake]. However, existing spoof detection systems are primarily formulated on reconstructed audio signals and are designed for post-hoc classification. They do not operate on discrete codec token sequences, nor are they intended to guide generation during decoding.
@@ 2.1 Overview
Pretrained codec-based language models have demonstrated strong capability in zero-shot speech synthesis by modeling discrete codec token sequences autoregressively. In this work, we adopt NeuTTS [^1] https://github.com/neuphonic/neutts , a pretrained codec-based TTS system, as the base generator, whose parameters remain fixed throughout our framework. While such models are trained to predict next tokens conditioned on ground-truth codec sequences, inference relies on previously generated tokens. This shift introduces a discrepancy between training and decoding, under which synthesized token sequences may progressively deviate from the natural codec distribution, as previously observed in [zhang2024speechalign].
@@ 2.2 Multi-Resolution Token-Level Spoof Detection
In addition, inspired by multi-scale discriminators in neural vocoders [kong2020hifi, leebigvgan], we introduce a resolution-based skip sampling strategy with downsampling rates $r\in\{1,2,5\}$. When $r=1$, the sequence remains at its native resolution, corresponding to the full 50-token span. Larger $r$ values subsample the sequence to form coarser token representations, encouraging the detector to uncover structural inconsistencies that may not be evident at the original granularity. While temporal cropping enhances coverage across time spans, skip sampling perturbs token resolution to probe structural regularity from a complementary perspective.
Each token segment is processed by an embedding layer followed by stacked Conformer [gulati2020conformer] blocks to jointly capture local correlations and long-range dependencies. The resulting sequence representation is aggregated via adaptive pooling and passed to a lightweight classifier to predict the probability of the segment being real or synthetic. The detector is trained with the binary cross-entropy (BCE) objective and optimized independently of the TTS backbone. The architecture is shared across all resolution settings, while parameters are learned separately for each configuration. Specifically, we train five models: $\mathcal{M}_{50}$, $\mathcal{M}_{25}$, and $\mathcal{M}_{10}$ for contiguous cropping with segment lengths $L=50,25,10$, respectively, and $\mathcal{M}_{50\leftarrow 25}$ and $\mathcal{M}_{50\leftarrow 10}$ for skip-sampled variants obtained by downsampling 50-token segments with rates $r=2$ and $r=5$, respectively.
@@ 2.3 Hierarchical Spoof-Guided Sampling
To improve the decoding robustness of autoregressive codec language models, we propose a hierarchical spoof-guided sampling framework that integrates Entropy-Aware Sampling (EAS) with the multi-resolution token-level spoof detection introduced in Section 2.2. We first adopt Entropy-Aware Sampling (EAS), shown in Algorithm 1, as the base decoding strategy. EAS is adapted from repetition-aware sampling (RAS) proposed in VALL-E 2 [chen2024vall], which penalizes recently generated tokens to mitigate looping behavior. However, conventional RAS relies on heuristic repetition counting and may over-suppress high-probability tokens, resulting in unstable decoding or excessive randomness. In contrast, EAS maintains a memory buffer that records competitive candidate tokens along with their rank positions and temporal age. The token-level penalty is modulated by inverse rank weighting and exponential temporal decay, with a clipping mechanism to prevent over-penalization. The adjusted distribution is then used for nucleus sampling, introducing entropy regularization while preserving distributional diversity.
@@ 3.1 Datasets
For spoof detection training, we use the LibriTTS [zen2019libritts] training split (approximately 100 hours of clean read English speech). LibriTTS is a multi-speaker corpus with careful segmentation and noise filtering. For each ground-truth utterance, we generate three synthetic counterparts using the same transcript but a different reference utterance from the same speaker under the default inference scheme. All real and synthetic waveforms are converted into discrete token sequences using the default speech tokenizer (NeuCodec).
For evaluation, experiments are conducted on both LibriSpeech [panayotov2015librispeech] and LibriTTS (test split), which provide diverse speakers and clean TTS-style speech, respectively. These datasets enable evaluation under standard speech conditions using WER, speaker similarity, and perceptual quality metrics. To further assess robustness under challenging phonetic patterns, we additionally evaluate on the TwistList benchmark [loakman-etal-2023-twistlist], which consists of linguistically constructed tongue twisters characterized by dense alliteration and repetitive phoneme patterns. We use the official test split and synthesize speech using the same conditioning and decoding pipeline as in the standard evaluation.
@@ 3.3 Evaluation Metrics
For objective evaluation, we report ASR-based intelligibility, speaker similarity, perceptual quality, and spoof detection performance. Intelligibility is measured using Word Error Rate (WER), computed with Whisper-large-v3 [radford2023robust] [^2] https://huggingface.co/openai/whisper-large-v3 for all decoding schemes. Speaker similarity is evaluated using the WavLM [chen2022wavlm] [^3] https://huggingface.co/microsoft/wavlm-base-plus-sv model by computing cosine similarity between L2-normalized embeddings of the generated utterance and another real utterance from the same speaker. Perceptual quality is assessed using neural MOS estimators, including NISQA [mittag2021nisqa] and MOSNet [lo2019mosnet], both reflecting overall speech quality.